### v1

# Beyond In-Domain Detection: SpikeScore for Cross-Domain Hallucination Detection

Abstract
Hallucination detection is critical for deploying large language models (LLMs) in real-world applications. Existing hallucination detection methods achieve strong performance when the training and test data come from the same domain, but they suffer from poor cross-domain generalization. In this paper, we study an important yet overlooked problem, termed generalizable hallucination detection (GHD), which aims to train hallucination detectors on data from a single domain while ensuring robust performance across diverse related domains. In studying GHD, we simulate multi-turn dialogues following LLMs initial response and observe an interesting phenomenon: hallucination-initiated multi-turn dialogues universally exhibit larger uncertainty fluctuations than factual ones across different domains. Based on the phenomenon, we propose a new score SpikeScore, which quantifies abrupt fluctuations in multi-turn dialogues. Through both theoretical analysis and empirical validation, we demonstrate that SpikeScore achieves strong cross-domain separability between hallucinated and non-hallucinated responses. Experiments across multiple LLMs and benchmarks demonstrate that the SpikeScore-based detection method outperforms representative baselines in cross-domain generalization and surpasses advanced generalization-oriented methods, verifying the effectiveness of our method in cross-domain hallucination detection.
1 Introduction
Hallucination detection (manakul-etal-2023-selfcheckgpt; farquhar2024detecting) is crucial for the reliable deployment of large language models (LLMs) in real-world applications, particularly in safety-critical domains such as education (harvey2025don), healthcare (roustan2025clinicians), and finance (kang2023deficiency). This is because LLMs may produce factually incorrect or logically inconsistent outputs, collectively referred to as hallucinations (10.1145/3703155), which can undermine user trust and lead to harmful consequences in high-stakes scenarios. To address this issue, hallucination detection has recently been widely studied (chang2024survey; minaee2024large).
Existing detection methods generally fall into two categories: training-free methods (janiak2025illusion; sun2025redeep) and training-based methods (obeso2025real; wei2024toward). Representative training-free methods (manakul-etal-2023-selfcheckgpt; farquhar2024detecting) rely on intrinsic model signals such as uncertainty, consistency, or attention patterns, while typical training-based methods (kossen2024semantic; wei2024toward) train lightweight feed-forward classifiers on hidden activations from specific layers to predict reliability. Training-based methods generally outperform training-free methods on in-domain test sets, and thus have become the mainstream direction.
Despite the notable progress achieved by training-based hallucination detection methods, these methods remain fundamentally limited by poor cross-domain generalization. Experiments in zhang2024prompt show that for training-based methods, such as SAPLMA (azaria2023internal) and SEP (kossen2024semantic), detection performance drops sharply when the test domain differs from the training domain. This domain sensitivity stems from reliance on domain-specific features learned during training, making them brittle to distribution shifts. Thus, despite dominating current research, training-based methods fail to provide generalizable detection for practical applications.
Figure 1: Case study of self-contradictions in multi-turn dialogues. SAPLMA scores (trained on Math, tested on CoQA—a representative cross-domain setup) with Llama-3.1-8B-Instruct reveal distinct trajectories: hallucination-initiated dialogues (left) display marked spikes in the uncertainty signal, in contrast to the relatively stable fluctuations of factual dialogues (right). Inspection of the dialogue content at spike locations shows that hallucinated responses oscillate between contradictory viewpoints and frequently reverse their stance (blue: prompts; red: errors; green: factual responses), whereas factual dialogues remain consistent with only minor self-corrections.
To mitigate the generalization issue in hallucination detection, we consider a more realistic yet challenging setting, termed generalizable hallucination detection (GHD), which aims to train hallucination detectors from a single domain while ensuring strong performance across diverse related domains. Prior work has also sought indicators that separate hallucinated from non-hallucinated responses (azaria2023internal; kulkarni2025evaluating; farquhar2024detecting; xu-etal-2023-understanding), but such studies mainly address challenge 1: ensuring sufficient separability between hallucinated and non-hallucinated responses (within a single domain). In contrast, GHD additionally requires challenge 2: maintaining this separability consistently across different domains.
An intriguing phenomenon provides crucial inspiration: LLMs often exhibit self-contradiction in multi-turn conversations (laban2025llms; tosato2025persistent), particularly when users shift their stance (ranaldi2023large; hong2025measuring), leading models to accommodate the new position and contradict their previous conclusions. We hypothesize that this behavior represents an explicit manifestation of internal uncertainty or confidence instability. When the entire dialogue stems from hallucinated responses, such self-contradictory or stance-shifting behaviors should occur more frequently, as they often trigger the model’s self-correction mechanism. Given the broad observation of this phenomenon, we are keen to explore its potential for identifying domain-invariant indicators.
To validate our intuition, we construct multi-turn dialogues by feeding the model’s initial response back as context for subsequent questions. Specifically, we compile a diverse prompt library (see Appendix G) containing follow-up questions that naturally extend from the initial query, simulating realistic conversational scenarios where each turn builds upon the previous response. Case studies (Figure 1) reveal that when initial responses are hallucinated, models produce self-contradictory replies across dialogue turns, causing SAPLMA (azaria2023internal) scores (a representative training-based method) to exhibit sharp spikes characterized by rapid rises followed by steep drops. In contrast, such spikes are notably absent in dialogues initiated from factual responses. We expect this phenomenon extends beyond individual cases, potentially providing a universal solution to the GHD problem and guiding further methodological exploration. This leads us to investigate:
Do hallucination-initiated multi-turn dialogues universally exhibit larger uncertainty fluctuations than factual ones across different domains?
We design SpikeScore, defined as the maximum second-order difference of SAPLMA scores along the induced continuation, to quantify these abrupt fluctuations in multi-turn dialogue paths. This metric captures the sharpness of peaks in the score trajectory by measuring the intensity of rapid rise-and-fall patterns, where larger values indicate dramatic confidence reversals. Through extensive empirical studies, we statistically validate the consistent patterns of SpikeScore across different datasets. Building on these statistical observations, we further establish Theorem 1, which proves that when the mean and variance satisfy certain constraints, SpikeScore addresses both challenges: it achieves a probabilistic lower bound for separability between hallucinated and non-hallucinated chains (challenge 1), and our analysis suggests that this bound may extend across domains under the examined conditions (challenge 2). In other words, our results indicate that SpikeScore can serve as an effective indicator for distinguishing hallucinated from factual responses, supported by statistical evidence and theoretical bound observed across multiple domains.
Extensive experiments are conducted on four LLMs (i.e., Llama-3.2-3B/3.1-8B, Qwen3-8B/14B), and six benchmarks (i.e., TriviaQA (joshi-etal-2017-triviaqa), CommonsenseQA (talmor-etal-2019-commonsenseqa), Belebele (costa-jussa-etal-2025-2m), CoQA (reddy-etal-2019-coqa), Math (hendrycksmath2021), and SVAMP (patel-etal-2021-nlp)) covering commonsense, knowledge-intensive, conversational, and mathematical reasoning. Empirical results in Section 4.2 demonstrate that the SpikeScore-based detection method consistently outperforms representative detection methods in cross-domain generalization and surpasses advanced generalization-oriented methods, e.g., PRISM (zhang2024prompt) and ICR Probe (zhang-etal-2025-icr). Our main contributions are summarized as follows:
$\bullet$ We exploit the intrinsic instability of LLMs in multi-turn dialogue by inducing post-answer continuations, yielding standardized trajectories for analysis.
$\bullet$ We introduce SpikeScore, a domain-invariant instability indicator that enables threshold-based hallucination detection, and theoretically prove a lower bound on its separability.
$\bullet$ Extensive experiments demonstrate that SpikeScore consistently achieves superior cross-domain performance compared to strong baselines across multiple models and datasets.
2 Learning Setups
LLMs and Token Sequences. Following NEURIPS2024_ba927059; oh2025understanding, we use a distribution $\mathbb{P}_{\bm{\theta}}(\cdot)$ over token sequences to define LLM, where $\bm{\theta}$ is the model parameters. Given a token sequence $\mathbf{Q}=[x_{1},\dots,x_{k}]$ representing the question, where each $x_{j}$ is the $j$-th token in the sequence. $\mathbb{P}_{\bm{\theta}}(\cdot)$ generates an answer $\mathbf{A}=[x_{k+1},\dots,x_{k+l}]$ by predicting each token based on the preceding context:
$\mathbb{P}_{\bm{\theta}}(x_{j}\mid x_{1},\dots,x_{j-1}),~\text{for }j=k+1,\dots,k+l.$ (1)
Domains and Datasets. Let $\mathcal{Q}$ and $\mathcal{A}$ be the spaces of questions and answers, respectively. We consider a ground-truth domain $\mathbb{P}_{Q,T}$, a joint distribution over $\mathcal{Q}\times\mathcal{A}$, where $Q$ and $T$ are the question and truthful-answer random variables, respectively.
Given $\mathbb{P}_{Q,T}$, each sample consists of a question $\mathbf{Q}$, a reference answer $\mathbf{A}_{\text{ref}}$, and optionally a context passage (if provided by the training dataset), where $\mathbf{A}_{\text{ref}}\sim\mathbb{P}_{T|Q}(\cdot\mid\mathbf{Q})$, here $\mathbb{P}_{T|Q}$ is the conditional distribution of $\mathbb{P}_{Q,T}$. For simplicity, we concatenate the context and the question into a single input sequence, which we denote as $\mathbf{Q}$. The dataset from $\mathbb{P}_{Q,T}$ can then be represented as $\mathcal{D}=\{(\mathbf{Q}_{1},\mathbf{A}_{\text{ref}_{1}}),\dots,(\mathbf{Q}_{n},\mathbf{A}_{\text{ref}_{n}})\}$, where $n$ is the size of samples.
Given a question $\mathbf{Q}\sim\mathbb{P}_{Q}$, the LLM $\mathbb{P}_{\bm{\theta}}(\cdot)$ generates an answer $\mathbf{A}$, i.e., $\mathbf{A}\sim\mathbb{P}_{\bm{\theta}}(\cdot|\mathbf{Q})$. Each generated answer $\mathbf{A}$ is then assigned a binary label $y\in\{0,1\}$ according to its semantic consistency with the reference answer $\mathbf{A}_{\text{ref}}$. Specifically, if $\mathbf{A}$ aligns with $\mathbf{A}_{\text{ref}}$, it is labeled as truthful (i.e., $y=0$); otherwise, it is labeled as hallucinated (i.e., $y=1$). The labeled dataset $\mathcal{D}_{l}$ is defined as:
$\mathcal{D}_{l}=\{(\mathbf{Q}_{1},\mathbf{A}_{1},y_{1}),\dots,(\mathbf{Q}_{n},\mathbf{A}_{n},y_{n})\}.$ (2)
Generalizable Hallucination Detection. In GHD, we consider a training domain $\mathbb{P}_{Q,T}$ and $N$ related test domains $\mathbb{P}_{Q,T}^{1},\ldots,\mathbb{P}_{Q,T}^{N}$. The goal of GHD is to train a detector on $\mathbb{P}_{Q,T}$ that generalizes well to the $N$ test domains. We formally introduce the definition of GHD below.
Problem 1 (Generalizable Hallucination Detection.)
Given a training dataset constructed from the training domain $\mathbb{P}_{Q,T}$ together with a given LLM $\mathbb{P}_{\bm{\theta}}$, i.e.,
$\mathcal{D}_{l}=\{(\mathbf{Q}_{1},\mathbf{A}_{1},y_{1}),\dots,(\mathbf{Q}_{n},\mathbf{A}_{n},y_{n})\},$ (3)
as introduced in Eq. (2), the objective of generalizable hallucination detection (GHD) is to learn a detector $D$ based on the LLM $\mathbb{P}_{\bm{\theta}}(\cdot)$ and the training dataset $\mathcal{D}_{l}$. The detector is required to generalize across $N$ test domains $\mathbb{P}_{Q,T}^{1},\ldots,\mathbb{P}_{Q,T}^{N}$, i.e., for any question–answer pair $(\mathbf{Q},\mathbf{A})$,
[…]
Note that when $N=1$ and $\mathbb{P}_{Q,T}^{1}=\mathbb{P}_{Q,T}$, GHD reduces to the classical hallucination detection.
Due to space constraints, the related work is discussed in Appendix LABEL:app:relwork.
3 Methodology
3.1 Exploring Cross-Domain Invariant Indicators
Recent studies (deshpande2025multichallenge; zhang2025survey; laban2025llms) have revealed a compelling phenomenon in multi-turn dialogues: when users repeatedly probe or challenge LLM responses, models often rapidly abandon their initial positions and defer to user suggestions, exhibiting marked self-contradiction. This behavior is remarkably consistent across diverse datasets, model families, and parameter scales, manifesting robustly regardless of conversational context.
This pattern is broadly consistent with evidence that self-contradictory behavior reflects underlying uncertainty in generated content (yoon2025reasoningmodelsbetterexpress; liu2024intrinsic). Building on this insight, we hypothesize that the degree of response instability in follow-up dialogues correlates with the factuality of the initial response. While all LLM responses may exhibit some level of self-contradiction when challenged, we expect hallucinated responses to demonstrate significantly higher instability. In contrast, factually grounded responses, while not immune to revision, should exhibit greater consistency when probed. Furthermore, given LLMs’ self-correction mechanisms (wang2024a; lee2025revise; zhao2025boosting; NEURIPS2024_5e5853f3), initially erroneous responses may trigger more dramatic shifts and sharper deviations as the model actively attempts to rectify its mistakes.
To validate our hypothesis, we construct a simulated dialogue environment to examine response dynamics across multiple turns. In our hallucination detection scenario, we treat the model’s initial response to a question as the starting point of the dialogue, with subsequent responses termed as continuation answers. The process of generating continuation answers is formulated as follows:
$\mathbf{A}^{k}\sim\mathbb{P}_{\bm{\theta}}(\cdot|\mathbf{Q},\mathbf{A}^{1},\mathbf{P}^{2},\mathbf{A}^{2},...,\mathbf{P}^{i},\mathbf{A}^{i},...,\mathbf{P}^{k}),~\text{for}~\text{any}~k>1,$ (5)
where $\mathbf{A}^{1}$ is the original answer $\mathbf{A}$, $\mathbf{A}^{i}$ is the $i$-th continuation answer ($i>1)$, and $\mathbf{P}^{i}$ is the $i$-th prompt to induce the generation of $i$-th continuation answer $\mathbf{A}^{i}$ for any $i>1$. The design of the prompt $\mathbf{P}^{i}$, motivated by chen2025learning and laban2025llms, is to induce contextually coherent responses given the preceding $i$ dialogue turns. This formulation enables us to simulate realistic multi-turn interactions. We provide the complete prompt specification in Appendix G.
Training Score for Uncertainity Estimation. To effectively evaluate the uncertainty of each continuation answer $\mathbf{A}^{k}$, we need to select suitable scoring methods. To further ensure that the scores remain relatively accurate in addressing hallucination detection, we adopt training-based methods that leverage information of training data tailored for hallucination detection.
In this work, we use the one of the most representative methods in hallucination detection, termed SAPLMA (azaria2023internal), which utilizes LLM’s internal state to reveal the truthfulness of a given answer. The details of SAPLMA is introduced as follows. Given the internal representation $\mathbf{E}_{\bm{\theta}}(\cdot)$ of the LLM $\mathbb{P}_{\bm{\theta}}(\cdot)$, we place a multilayer perceptron (MLP) followed by a sigmoid activation on top of $\mathbf{E}_{\bm{\theta}}(\cdot)$ to produce a probabilistic output $p_{\mathbf{W}}(\cdot)\in[0,1]$. The parameters $\mathbf{W}$ are then optimized using the cross-entropy loss over the training data $\mathcal{D}_{l}=\{(\mathbf{Q}_{i},\mathbf{A}_{i},y_{i})\}_{i=1}^{n}$, i.e.,
$\begin{split}\widehat{\mathbf{W}}\in\operatorname*{arg\,min}_{\mathbf{W}}\mathcal{L}(\mathbf{W};\mathcal{D}_{l})=-\frac{1}{n}\sum_{i=1}^{n}\Big(&y_{i}\log p_{\mathbf{W}}(\mathbf{E}_{\bm{\theta}}(\mathbf{A}_{i}|\mathbf{Q}_{i}))\\ +&(1-y_{i})\log\big(1-p_{\mathbf{W}}(\mathbf{E}_{\bm{\theta}}(\mathbf{A}_{i}|\mathbf{Q}_{i}))\big)\Big).\end{split}$ (6)
Next, to evaluate the uncertainty of a continuation answer $\mathbf{A}^{k}$, we use the following formulation:
$S(\mathbf{A}^{k};\mathbf{Q},\mathbf{A})=p_{\widehat{\mathbf{W}}}(\mathbf{E}_{\bm{\theta}}(\mathbf{A}^{k}|\mathbf{Q},\mathbf{A}^{1},\mathbf{P}^{2},\mathbf{A}^{2},...,\mathbf{P}^{i},\mathbf{A}^{i},...,\mathbf{P}^{k})).$ (7)
Finally, for each $\mathbf{Q}$ and $\mathbf{A}$, we obtain a score sequence:
$\mathbf{S}(\mathbf{Q},\mathbf{A})=[S(\mathbf{A}^{1};\mathbf{Q},\mathbf{A}),S(\mathbf{A}^{2};\mathbf{Q},\mathbf{A}),...,S^{K}(\mathbf{A}^{1};\mathbf{Q},\mathbf{A})],$ (8)
where $K$ is the stoping step. In our experiment, we set $K=20$ as the defult.
Remark. We also evaluate other representative training-based scoring method SEP (kossen2024semantic), as well as several training-free methods, including Perplexity (10.1145/3571730), Reasoning score (sun2025detection), and In-Context Sharpness (10.5555/3692070.3692364). These methods yield conclusions consistent with those of SAPLMA. Moreover, we observe that integrating our SpikeScore with SAPLMA or SEP achieves substantially better overall performance than training-free methods, showing the clear advantage of training-based methods. Please see Section 4.2.
SpikeScore: Estimating Score Sequence Fluctuations. There are many methods to evaluate the sequence fluctuations, such as computing the variance or extreme gap. It is worth noting that sun2025detection proposes the coefficient of variation score, which measures the score sequence fluctuations using the coefficient of variation, i.e., the standard deviation normalized by the expection.
The coefficient of variation score primarily captures the overall fluctuation of a sequence. However, focusing solely on global variation may overlook fine-grained differences (9722882). Measuring local fluctuations offers a more sensitive and discriminative characterization of sequence fluctuations (jennings2004variance). Motivated by qian2025structural, we propose to measure local fluctuations using the maximum second-order difference, i.e., given a score sequence $\mathbf{S}(\mathbf{Q},\mathbf{A})$,
$\text{Max}|{\Delta}^{2}|(\mathbf{S}(\mathbf{Q},\mathbf{A}))=\max_{1<k<K-1}|S(\mathbf{A}^{k+1};\mathbf{Q},\mathbf{A})-2S(\mathbf{A}^{k};\mathbf{Q},\mathbf{A})+S(\mathbf{A}^{k-1};\mathbf{Q},\mathbf{A})|.$ (9)
$\text{Max}|{\Delta}^{2}|$ captures the maximum curvature of the score sequence $\mathbf{S}(\mathbf{Q},\mathbf{A})$, corresponding to the most pronounced local fluctuation. In other words, the maximum second-order difference quantifies the point of greatest irregularity or instability within the score sequence $\mathbf{S}(\mathbf{Q},\mathbf{A})$. In this work, we call the maximum second-order difference $\text{Max}|{\Delta}^{2}|$ as SpikeScore.
Remark. We further compare SpikeScore with the coefficient of variation score. Empirically, SpikeScore outperforms the coefficient of variation score. Details are provided in Appendix LABEL:app:addexp_spike_vs_cv.
3.2 Evaluations of SpikeScore
In this part, we conduct experiments to examine the invariance of SpikeScore across hallucinated and non-hallucinated answers in different domains. We use six datasets: TriviaQA (joshi-etal-2017-triviaqa), CommonsenseQA (talmor-etal-2019-commonsenseqa), Belebele (costa-jussa-etal-2025-2m), CoQA (reddy-etal-2019-coqa), Math (hendrycksmath2021), and SVAMP (patel-etal-2021-nlp). In each group of experiments, one dataset is used for training and the remaining five are used for testing, yielding six groups of experiments in total (see Appendix LABEL:app:details for details). The experimental results are summarized in Figures 2 and 3, from which we highlight key observations.
Observation 1 (Expectation Invariance.)
Let $\mathbb{P}_{Q,T}^{t}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{P}_{Q,T}^{i}$ be the non-hallucination domain as the uniform mixture of $N$ test domains $\mathbb{P}_{Q,T}^{1},...,\mathbb{P}_{Q,T}^{N}$ introduced in Section 2. The value of SpikeScore trained on the associated training domain $\mathbb{P}_{Q,T}$ satisfies that
$2\mathbb{E}_{(\mathbf{Q},\mathbf{A})\sim\mathbb{P}_{Q,T}^{t}}\textbf{\rm Max}|{\Delta}^{2}|(\mathbf{S}(\mathbf{Q},\mathbf{A}))<\mathbb{E}_{(\mathbf{Q},\mathbf{H})\sim\mathbb{P}_{Q,H}^{t}}\textbf{\rm Max}|{\Delta}^{2}|(\mathbf{S}(\mathbf{Q},\mathbf{H})),$ (10)
where $\mathbb{P}_{Q,H}^{t}$ is the hallucination domain corresponding to $\mathbb{P}_{Q,T}^{t}$, the marginal distributions related to the question random variable $Q$ of $\mathbb{P}_{Q,T}^{t}$ and $\mathbb{P}_{Q,H}^{t}$ are consistent.
Based on the comparison in Figure 2 (a), we find that the expectation of SpikeScore for the hallucination domain is more than twice that for the non-hallucination domain. This result reflects the separability of hallucination and non-hallucination domains captured by SpikeScore. Although this result aligns with our expectation, it is based solely on the expected SpikeScore. If SpikeScore exhibits large variance, the separability between the hallucination and non-hallucination domains remains questionable. Hence, we further study the standard deviation of SpikeScore in Figure 2 (b).
(a) Mean values across training datasets
(b) Standard deviations across training datasets
Figure 2: Statistical properties of SpikeScore for hallucination detection across cross-domain scenarios. Using Llama-3.1-8B-Instruct, we train uncertainty probes on individual datasets (e.g., CoQA) and evaluate on an equal mixture of the remaining datasets (TriviaQA, CommonsenseQA, Belebele, Math, SVAMP). Figure (a) demonstrates that hallucinated dialogues produce higher mean SpikeScore values compared to factual dialogues across all training configurations, indicating robust discriminative capability. Figure (b) shows that hallucinated dialogues exhibit noticeably higher standard deviations compared to factual ones. Although the variance gap is larger, Theorem 1 guarantees that separability between hallucinated and factual domains still holds under controlled coefficient-of-variation conditions.
Observation 2 (Standard Deviation Differences.)
For the non-hallucination domain $\mathbb{P}_{Q,T}^{t}$ introduced in Observation 1, the value of SpikeScore trained on the associated training domain $\mathbb{P}_{Q,T}$ satisfies that
$1<\frac{\mathbf{Std}_{(\mathbf{Q},\mathbf{H})\sim\mathbb{P}_{Q,H}^{t}}\textbf{\rm Max}|{\Delta}^{2}|(\mathbf{S}(\mathbf{Q},\mathbf{H}))}{\mathbf{Std}_{(\mathbf{Q},\mathbf{A})\sim\mathbb{P}_{Q,T}^{t}}\textbf{\rm Max}|{\Delta}^{2}|(\mathbf{S}(\mathbf{Q},\mathbf{A}))}\leq 2.5,$
where $\mathbb{P}_{Q,H}^{t}$ is the hallucination domain introduced in Observation 1.
In Figure 2 (b), we observe that the standard deviation of SpikeScore in the hallucination domain is larger than in the non-hallucination domain, with the ratio of standard deviations reaching $2.5$, which exceeds the expectation ratio of $2$ reported in Observation 1. This indicates that variance differences alone may hinder direct separability. Nevertheless, as shown in Theorem 1, separability between hallucination and non-hallucination domains can still be established under controlled coefficient-of-variation conditions (i.e., variance normalized by expectation). See Theorem 1 for details.
Theorem 1
Suppose Observations 1 and 2 hold. If the coefficient of variation for $\mathbb{P}_{Q,T}^{t}$ satisfies
$\mathbf{CV}_{(\mathbf{Q},\mathbf{A})\sim\mathbb{P}_{Q,T}^{t}}\textbf{\rm Max}|{\Delta}^{2}|(\mathbf{S}(\mathbf{Q},\mathbf{A}))=\frac{\mathbf{Std}_{(\mathbf{Q},\mathbf{A})\sim\mathbb{P}_{Q,T}^{t}}\textbf{\rm Max}|{\Delta}^{2}|(\mathbf{S}(\mathbf{Q},\mathbf{A}))}{\mathbb{E}_{(\mathbf{Q},\mathbf{A})\sim\mathbb{P}_{Q,T}^{t}}\textbf{\rm Max}|{\Delta}^{2}|(\mathbf{S}(\mathbf{Q},\mathbf{A}))}\leq 0.1\cdot t,$
for some $t>0$, then we have
$\mathbb{P}\big(\textbf{\rm Max}|{\Delta}^{2}|(\mathbf{S}(\mathbf{Q}^{\prime},\mathbf{H}^{\prime}))>\textbf{\rm Max}|{\Delta}^{2}|(\mathbf{S}(\mathbf{Q},\mathbf{A})\big)\geq\frac{1}{1+0.0725\cdot t^{2}},$
where $(\mathbf{Q}^{\prime},\mathbf{H}^{\prime})\sim\mathbb{P}_{Q,H}^{t}$ is the hallucinated sample and $(\mathbf{Q},\mathbf{A})\sim\mathbb{P}_{Q,T}^{t}$ is the factual sample.
Proof. The proof of Theorem 1 can be found in Appendix LABEL:app:proofs.
Theorem 1 states that under Observations 1 and 2, if the coefficient of variation of SpikeScore in the hallucination domain is sufficiently small ($0.1\cdot t$ in Theorem 1), then SpikeScore in the hallucination domain will, with high probability, exceed that in the non-hallucination domain. Therefore, we further study the coefficient of variation of SpikeScore through experiments.
In Figure 3, we report the coefficient of variation of SpikeScore across six groups of experiments, conducted under the same settings as in Figure 2. From this, we derive the following observation.
Observation 3 (Controlled Coefficient of Variation.)
For the non-hallucination domain $\mathbb{P}_{Q,T}^{t}$, the value of SpikeScore trained on the training domain $\mathbb{P}_{Q,T}$ satisfies that
$\mathbf{CV}_{(\mathbf{Q},\mathbf{A})\sim\mathbb{P}_{Q,T}^{t}}\textbf{\rm Max}|{\Delta}^{2}|(\mathbf{S}(\mathbf{Q},\mathbf{A}))\leq 0.2.$ (11)
Figure 3 shows that the coefficient of variation for non-hallucination domain does not exceed $0.2$. Then, Theorem 1 implies that the SpikeScore in the hallucination domain will exceed that in the non-hallucination domain with probability at least $0.775$, i.e.,
$\mathbb{P}\big(\textbf{\rm Max}|{\Delta}^{2}|(\mathbf{S}(\mathbf{Q},\mathbf{H}))>\textbf{\rm Max}|{\Delta}^{2}|(\mathbf{S}(\mathbf{Q},\mathbf{A})\big)\geq\frac{1}{1+0.0725\cdot 2^{2}}\approx 0.775.$
Figure 3: The coefficient of variation of SpikeScore based on non-hallucinated examples across datasets. Bars represent the coefficient of variation calculated from the maximum second-order difference within the first 20 steps. The experimental setup is identical to Figure 2. All coefficient of variation values remain below 0.2.
We defer a fuller discussion of separability to Appendix E, where we also compare SpikeScore with alternative indicators across models and observe generally stronger and more consistent performance for SpikeScore across a range of evaluation settings and representative cross-domain scenarios.
Through the above analyses, we have shown that SpikeScore is able to achieve good separability between hallucination and non-hallucination domains with high probability. It should be noted that, to ensure robust cross-domain evaluation, we follow a commonly adopted practice in cross-domain generalization studies (hendrycks2018deep; 10.5555/3495724.3497526; 9879414; 10.5555/3737916.3739021). Specifically, after training on one dataset, we evaluate on a held-out pool formed by uniformly mixing the remaining datasets, a protocol commonly used to measure cross-domain generalization. From a probabilistic perspective, this pooled evaluation is equivalent to taking the expectation of separability over individual test domains, and thus faithfully reflects cross-domain generalization under diverse conditions.
3.3 Leveraging SpikeScore for GHD
In this part, we leverage SpikeScore to construct our method for GHD. Given a question $\mathbf{Q}$ and the corresponding answer $\mathbf{A}$, we first compute the SpikeScore of $\mathbf{A}$ according to Eq. (9). Based on this score, we then apply the following formulation to determine whether $\mathbf{A}$ is hallucinated or non-hallucinated: given a thershold $\lambda>0$,
$D_{\lambda}(\mathbf{Q},\mathbf{A})=\begin{cases}0,&\text{if }\text{Max}\left|\Delta^{2}\right|(\mathbf{S}(\mathbf{Q},\mathbf{A}))<\lambda,\\ 1,&\text{otherwise},\end{cases}$ (12)
where $1$ indicates that $\mathbf{A}$ is hallucinated, and 0 indicates that $\mathbf{A}$ is non-hallucinated.
Table 1: Cross-domain hallucination detection performance. To ensure fair comparison, we use a leave-one-out protocol: training-based methods train on each dataset (columns) while all methods are evaluated on the remaining five datasets, preventing training-based methods from being tested on their own training domain (see Appendix LABEL:appendix:raw_results_explanation for details). The table is split into two blocks only due to space constraints, corresponding to different model families. We report mean AUROC across test datasets. Best results are bold, second-best are underlined.
Method Training Dataset
TriviaQA CommonsenseQA Belebele CoQA Math SVAMP Average
[…]
Llama Llama Llama Llama Llama Llama Llama Llama Llama Llama Llama Llama Llama Llama
3.2-3B 3.1-8B 3.2-3B 3.1-8B 3.2-3B 3.1-8B 3.2-3B 3.1-8B 3.2-3B 3.1-8B 3.2-3B 3.1-8B 3.2-3B 3.1-8B
[…]
Perplexity 0.5876 0.6353 0.5893 0.6379 0.5995 0.6445 0.5962 0.6420 0.5976 0.6448 0.6013 0.6505 0.5953 0.6425
Semantic Entropy 0.6123 0.6615 0.6250 0.6725 0.6320 0.6769 0.6133 0.6615 0.6307 0.6812 0.6149 0.6666 0.6214 0.6700
EigenScore 0.6431 0.6945 0.6378 0.6895 0.6515 0.7038 0.6417 0.6876 0.6515 0.7043 0.6503 0.6979 0.6460 0.6963
Lexical Similarity 0.6299 0.6961 0.6283 0.6911 0.6305 0.6882 0.6254 0.6790 0.6333 0.6946 0.6379 0.7014 0.6309 0.6917
Verbalize 0.5045 0.5520 0.4990 0.5476 0.5077 0.5560 0.4873 0.5400 0.4965 0.5436 0.4931 0.5420 0.4980 0.5469
InterrogateLLM 0.6508 0.6992 0.6528 0.7030 0.6449 0.6993 0.6456 0.7013 0.6531 0.7041 0.6544 0.7050 0.6503 0.7020
[…]
$M$ 0.5787 0.5784 0.5468 0.5384 0.4800 0.4893 0.5948 0.6120 0.5767 0.5841 0.5778 0.5788 0.5591 0.5635
SEP 0.5597 0.5417 0.5170 0.5102 0.5358 0.5004 0.4984 0.5221 0.5390 0.5420 0.5119 0.5529 0.5270 0.5282
SAPLMA 0.5723 0.6141 0.5319 0.5217 0.4933 0.5074 0.6468 0.6358 0.5861 0.6003 0.5855 0.5790 0.5693 0.5764
[…]
PRISM 0.6817 0.6971 0.6988 0.7051 0.7067 0.6974 0.7984 0.7917 0.6051 0.6475 0.6811 0.6784 0.6953 0.7029
ICR Probe 0.7464 0.7310 0.7118 0.7205 0.7416 0.7498 0.8155 0.8074 0.7615 0.7487 0.7010 0.7058 0.7463 0.7439
[…]
Qwen3 Qwen3 Qwen3 Qwen3 Qwen3 Qwen3 Qwen3 Qwen3 Qwen3 Qwen3 Qwen3 Qwen3 Qwen3 Qwen3
-8B -14B -8B -14B -8B -14B -8B -14B -8B -14B -8B -14B -8B -14B
[…]
Perplexity 0.5963 0.6222 0.6108 0.6287 0.6153 0.6311 0.6090 0.6284 0.6126 0.6343 0.6224 0.6362 0.6111 0.6302
Semantic Entropy 0.6295 0.6499 0.6385 0.6664 0.6368 0.6684 0.6302 0.6532 0.6481 0.6785 0.6310 0.6569 0.6357 0.6622
EigenScore 0.6651 0.6942 0.6560 0.6864 0.6727 0.7055 0.6538 0.6809 0.6753 0.6995 0.6636 0.6978 0.6644 0.6940
Lexical Similarity 0.6491 0.6683 0.6402 0.6626 0.6423 0.6632 0.6397 0.6571 0.6422 0.6625 0.6493 0.6682 0.6438 0.6636
Verbalize 0.5273 0.5402 0.5232 0.5226 0.5228 0.5345 0.5172 0.5137 0.5090 0.5308 0.5123 0.5237 0.5186 0.5276
InterrogateLLM 0.6662 0.6845 0.6670 0.6814 0.6611 0.6777 0.6761 0.6787 0.6771 0.6852 0.6718 0.6904 0.6699 0.6830
[…]
$M$ 0.5858 0.5755 0.5309 0.5411 0.4866 0.4854 0.6161 0.6037 0.5789 0.5958 0.5808 0.5795 0.5632 0.5635
SEP 0.5395 0.5273 0.5157 0.5164 0.4923 0.4952 0.5164 0.5275 0.5309 0.5550 0.5326 0.5430 0.5212 0.5274
SAPLMA 0.6344 0.6108 0.5163 0.5281 0.4985 0.5119 0.6301 0.6392 0.5685 0.6063 0.5753 0.5760 0.5705 0.5787
[…]
PRISM 0.7017 0.6923 0.7075 0.7187 0.6887 0.7097 0.7891 0.8116 0.6516 0.6427 0.6809 0.6681 0.7032 0.7072
ICR Probe 0.7317 0.7337 0.7249 0.7186 0.7271 0.7488 0.8000 0.8084 0.7281 0.7434 0.7170 0.7078 0.7381 0.7435
[…]
4 Experiment
4.1 Experimental Setups
We evaluate SpikeScore on six widely used hallucination detection benchmarks: CommonsenseQA (talmor-etal-2019-commonsenseqa), TriviaQA (joshi-etal-2017-triviaqa), Belebele (costa-jussa-etal-2025-2m), CoQA (reddy-etal-2019-coqa), Math (hendrycks2021measuring), and SVAMP (patel-etal-2021-nlp). Baselines include Perplexity (ren2023outofdistribution), Semantic Entropy (farquhar2024detecting), EigenScore (chen2024inside), Lexical Similarity (lin2024generating), Verbalize (lin2022teaching), InterrogateLLM (yehuda-etal-2024-interrogatellm), MM (marks2023geometry), SEP (kossen2024semantic; farquhar2024detecting), SAPLMA (azaria2023internal), PRISM (SAPLMA instantiation) (zhang2024prompt), and ICR Probe (zhang-etal-2025-icr), with AUROC as the primary evaluation metric. We test across four carefully selected open-source LLMs—Llama-3.2-3B-Instruct and Llama-3.1-8B-Instruct (abs-2407-21783), and Qwen3-8B-Instruct and Qwen3-14B-Instruct (qwen3)—to cover diverse model families and parameter scales. Additional dataset sampling strategies, baseline configurations, and hyperparameters are provided in detail in Appendix LABEL:app:details.
4.2 Experimental Results
Overall Cross-domain Hallucination Detection Performance. In Table 3.3, we compare the proposed SpikeScore with three categories of baselines: training-free methods that avoid generalization issues, traditional training-based methods, and cross-domain specialized approaches. SpikeScore consistently achieves the highest average AUROC across all model families and maintains superior or competitive performance in individual dataset evaluations. Traditional training-based methods (MM, SEP, SAPLMA) generally exhibit poor cross-domain performance, while training-free methods show reasonable adaptability but fall short of specialized cross-domain approaches like PRISM and ICR Probe. Notably, SpikeScore demonstrates improved performance with larger models within the same family, suggesting that enhanced self-correction capabilities in larger LLMs (yang-etal-2025-confidence) generate more detectable attention spike patterns during conversational refinement.
Compatibility with Different Backbone Scoring Methods. To validate the generalizability of our spike-based detection framework, we comprehensively evaluate SpikeScore with various backbone scoring methods, including both training-free approaches and training-based methods. As shown in Table 1, all combinations demonstrate effective hallucination detection performance across different LLMs and datasets, strongly suggesting that the spike phenomenon we identified is a general and transferable characteristic of hallucinated outputs rather than specific to particular scoring mechanisms. Moreover, training-based methods (SpikeScore+SEP and SpikeScore+SAPLMA) achieve substantially better performance than their training-free counterparts, further confirming the practical advantage of learning-based approaches in cross-domain hallucination detection. Implementation details for adapting these scoring methods to our framework are provided in Appendix LABEL:app:details_backbones.
Extended Analyses and Ablations. We conduct extension studies detailed in Appendix LABEL:app:addexp. First, we analyze cross-domain generalization (Appendix LABEL:app:addexp_why_Robustness): SpikeScore captures low-frequency, domain-agnostic instability patterns through multi-turn dialogue dynamics, enabling rapid convergence with minimal training data, whereas single-step methods rely on high-frequency, semantically-entangled features that fail to transfer. Second, train-test heatmaps (Appendix LABEL:app:addexp_heatmaps) confirm strong in-domain performance with relatively clear diagonal patterns, though cross-domain evaluation remains our primary focus. Third, ablations against coefficient of variation (Appendix LABEL:app:addexp_spike_vs_cv) demonstrate that second-order differences consistently outperform first-order variability measures, validating our curvature-based approach. Finally, dialogue length analysis (Appendix LABEL:app:addexp_steps20) shows that performance saturates around 15–20 turns, which motivates the default setting of K=20K{=}20. Based on this finding, we regard the first 20 dialogue turns as the most informative region for hallucination detection and refer to them as the early stage throughout this paper, rather than extending dialogues indefinitely.
Table 2: Performance of SpikeScore combined with different backbone scoring methods. Following the leave-one-out protocol from Table 3.3, we evaluate SpikeScore integrated with both training-free methods (Perplexity, Reasoning score, In-Context Sharpness) and training-based methods (SEP, SAPLMA) across six datasets using different LLMs. The table is split into two blocks only due to space constraints, corresponding to different model families. We report mean AUROC across test datasets. Best results are bold, second-best are underlined.
Method Leave-one-out Dataset
TriviaQA CommonsenseQA Belebele CoQA Math SVAMP Average
[…]
Llama Llama Llama Llama Llama Llama Llama Llama Llama Llama Llama Llama Llama Llama
3.2-3B 3.1-8B 3.2-3B 3.1-8B 3.2-3B 3.1-8B 3.2-3B 3.1-8B 3.2-3B 3.1-8B 3.2-3B 3.1-8B 3.2-3B 3.1-8B
[…]
SpikeScore+Perplexity 0.6541 0.6942 0.6254 0.6755 0.6505 0.7033 0.8094 0.7936 0.6978 0.7225 0.6357 0.7066 0.6788 0.7160
SpikeScore+Reasoning score 0.7208 0.7282 0.6559 0.7211 0.6924 0.7504 0.8222 0.8267 0.7471 0.7794 0.6795 0.7514 0.7196 0.7595
SpikeScore+In-Context Sharpness 0.6772 0.7334 0.6615 0.6867 0.6505 0.7098 0.8185 0.7816 0.7170 0.7498 0.6836 0.7250 0.7014 0.7311
[…]
SpikeScore+SEP 0.7230 0.7389 0.6824 0.7341 0.6873 0.7518 0.8315 0.8405 0.7470 0.7888 0.7051 0.7565 0.7294 0.7684
[…]
Qwen3 Qwen3 Qwen3 Qwen3 Qwen3 Qwen3 Qwen3 Qwen3 Qwen3 Qwen3 Qwen3 Qwen3 Qwen3 Qwen3
-8B -14B -8B -14B -8B -14B -8B -14B -8B -14B -8B -14B -8B -14B
[…]
SpikeScore+Perplexity 0.6717 0.7051 0.6693 0.6692 0.6112 0.6926 0.7458 0.7382 0.6870 0.7533 0.6598 0.7096 0.6742 0.7113
SpikeScore+Reasoning score 0.7055 0.7118 0.7018 0.7179 0.6650 0.7012 0.7955 0.7952 0.7305 0.7750 0.6995 0.7314 0.7163 0.7387
SpikeScore+In-Context Sharpness 0.6692 0.7128 0.7095 0.7219 0.6574 0.7174 0.7923 0.7779 0.6869 0.7566 0.6899 0.7434 0.7008 0.7383
[…]
SpikeScore+SEP 0.7190 0.7548 0.7172 0.7451 0.7021 0.7484 0.7972 0.8244 0.7346 0.7918 0.7060 0.7661 0.7293 0.7718
[…]
4.3 Extending SpikeScore to Retrieval-Augmented Settings
SpikeScore is a post-hoc framework that operates entirely on the multi-turn continuation trajectories after an initial answer is produced. This property makes it naturally compatible with structured pipelines such as retrieval-augmented generation (RAG): regardless of whether the upstream system performs retrieval, tool calling, or other pipeline operations, once a natural language answer is available, we can induce self-dialogue, extract turn-level hallucination scores, and compute SpikeScore over the resulting sequence.
To examine this in realistic RAG scenarios, we build a retrieval pipeline using TriviaQA (joshi-etal-2017-triviaqa) and RAGTruth (niu-etal-2024-ragtruth), two widely used benchmarks in RAG research. Rather than relying on the datasets’ predefined question–context pairs, we treat all evidence passages as a shared corpus, embed queries and documents with bge-large-en (bge_embedding), and retrieve the top-4 candidates via a Faiss (johnson2019billion) index. The retrieved passages are concatenated with the question and fed to the same LLMs used in the main dialogue experiments. Importantly, all training-based methods—including SpikeScore—are trained only on non-RAG CoQA, so evaluation on TriviaQA and RAGTruth becomes a strict cross-domain generalization test. Detailed pipeline construction is provided in Appendix LABEL:app:rag.
Table LABEL:tab:rag_results_main reports the AUROC on both RAG datasets. Training-free baselines degrade substantially under retrieval-induced noise, and training-based methods also lose accuracy when transferred from dialogue-style data to retrieval-based pipelines. In contrast, SpikeScore consistently outperforms all baselines, including the strongest competitor (ICR Probe), across every model scale and on both datasets. This robustness indicates that SpikeScore’s curvature-based temporal features remain discriminative even when hallucinations stem from imperfect retrieval or incomplete evidence.
We further observe that TriviaQA and RAGTruth exhibit highly aligned trends: SpikeScore remains the leading method across different LLM families and parameter sizes, and its performance margins over baselines are stable across datasets. Together with CoQA—which corresponds to an “idealized RAG” scenario where retrieval is perfect—these results show that SpikeScore generalizes reliably from open-ended dialogue to structured, tool-augmented workflows. This suggests that the temporal curvature features captured by SpikeScore are not tied to any specific interaction pattern but instead reflect a deeper regularity in how models behave when drifting toward hallucination. As retrieval-augmented systems become increasingly central in agentic architectures, such robustness is especially valuable, indicating that SpikeScore may serve as a unifying post-hoc indicator across both free-form and pipeline-driven LLM deployments.
Qwen-2.5-7B-ConsistentChat is an instruction-tuned model derived from the Qwen-2.5-7B base. It is trained on the ConsistentChat dataset to specifically mitigate topic drift and improve dialogue coherence.
  • Qwen3-4B-Instruct vs. EvolLLM-Linh. EvolLLM-Linh is fine-tuned from Qwen3-4B-Instruct-2507. While primarily designed to enhance robustness and accuracy in tool-using scenarios, it is optimized for multi-turn coherence. We evaluate it in a standard chat setting (without tools) to assess its inherent consistency capabilities relative to the base model.
    In both pairs, the base checkpoints serve as controls that have not been explicitly optimized for multi-turn consistency beyond standard instruction tuning.
  • Experimental Setup.
    We adhere to the identical GHD protocol used in the main experiments:
    • Datasets: The same six benchmarks (TriviaQA, CommonsenseQA, Belebele, CoQA, Math, and SVAMP);
    • Protocol: Identical prompts, generation parameters, and dialogue lengths;
    • Metric: The same SpikeScore computation and AUROC evaluation.
    The only variable in this ablation is the underlying model checkpoint.
    Results.
    Table 8 reports the AUROC for each dataset and the average across all domains.
    Table 8: Effect of multi-turn consistency training on SpikeScore performance. We compare base models with their counterparts explicitly optimized for multi-turn consistency. Values denote AUROC.
    Model TriviaQA CSQA Belebele CoQA Math SVAMP Avg.
    Qwen-2.5-7B 0.7242 0.7269 0.6928 0.8257 0.7521 0.7265 0.7414
    Qwen-2.5-7B-ConsistentChat 0.7560 0.7463 0.6879 0.8148 0.7295 0.7419 0.7461
    Qwen3-4B-Instruct 0.7357 0.7367 0.6774 0.8505 0.7484 0.7335 0.7470
    EvolLLM-Linh 0.7430 0.7497 0.7159 0.8213 0.7531 0.7085 0.7486
    As shown in Table 8, the AUROC scores remain highly stable—and even slightly improved on average—after multi-turn consistency training. For instance, the average AUROC for Qwen-2.5-7B-ConsistentChat increases to 0.7461 from 0.7414.
    Crucially, we observe no collapse in separability. This indicates that the curvature irregularities detected by SpikeScore are not artifacts of training inconsistency, but rather robust indicators of hallucination that persist even in models optimized for smooth multi-turn interactions.
    Discussion.
    These results are consistent with the following intuition. Multi-turn consistency training tends to smooth and stabilize the trajectories for both hallucinated and non-hallucinated cases simultaneously: it encourages the model to respond more consistently overall, but does not selectively “flatten out” only the spike-like behaviour associated with hallucinations. In other words, the consistency of hallucinated and non-hallucinated trajectories increases in roughly parallel fashion. Because SpikeScore is always computed within a single model to compare hallucinated versus factual trajectories, such parallel shifts do not eliminate the relative curvature gap between the two classes.
    This ablation therefore supports our claim that SpikeScore behaves as a robust hallucination-detection signal, rather than merely reflecting the strength of a particular multi-turn consistency training pipeline. The spike-like curvature it exploits appears to be an independent, cross-domain stable signature of hallucination, even in models that have been explicitly strengthened for multi-turn coherence.
    Appendix E Discussion on Separability
    Table 9: Separability (mixture-domain lower bound) across models and methods. Best results are bold, second-best are underlined.
    Method Llama Models Qwen Models
    Llama3.2-3B Llama3.1-8B Qwen3-8B Qwen3-14B
[…]
    $M$ 0.538 0.545 0.509 0.527
    SEP 0.503 0.507 0.512 0.511
    SAPLMA 0.516 0.541 0.569 0.563
[…]
    PRISM 0.647 0.670 0.633 0.672
    ICR Probe 0.695 0.702 0.686 0.705
[…]
    Separability, as introduced in Theorem 1, characterizes the probability that hallucinated responses achieve higher SpikeScore values than factual ones under a mixture-domain evaluation. Specifically, instead of evaluating on each domain separately, we form a mixed test pool across domains and compute
    $\mathrm{AUROC}_{\mathrm{mix}}=\mathbb{P}(X>Y),\quad X\sim\mathcal{H},~Y\sim\mathcal{A},$
    where $\mathcal{H}$ and $\mathcal{A}$ denote the hallucination and non-hallucination distributions in the mixture. Our theoretical analysis provides a distribution-free lower bound on this probability, i.e., a guaranteed separability level, even under worst-case variance conditions. Importantly, this separability lower bound is not the same as the per-domain AUROCs reported in the main text. The AUROC values in Table 3.3 are obtained by training on one dataset and testing on each remaining dataset separately, followed by averaging. In contrast, the theoretical separability bound aggregates cross-domain pairs into a single mixed evaluation. Because the evaluation protocols differ, it is possible—indeed expected—that some individual domain AUROCs fall below the mixture-level lower bound, even though the average AUROC remains higher.
    Formally, for a single domain $A$ with hallucination and non-hallucination distributions $(X_{A},Y_{A})$, Cantelli’s inequality yields
    $\mathrm{AUROC}_{A}\;\geq\;\frac{(\delta_{A}-1)^{2}}{(r_{A}^{2}+1)c_{A}^{2}+(\delta_{A}-1)^{2}},$
    where $\delta_{A}=\tfrac{\mathbb{E}X_{A}}{\mathbb{E}Y_{A}}$ is the mean ratio, $r_{A}=\tfrac{\mathrm{Std}(X_{A})}{\mathrm{Std}(Y_{A})}$ is the variance ratio, and $c_{A}=\tfrac{\mathrm{Std}(Y_{A})}{\mathbb{E}Y_{A}}$ is the coefficient of variation of non-hallucinations. If domain $A$ happens to exhibit weaker mean gaps ($\delta_{A}\!\approx\!1$), larger variance inflation ($r_{A}$ large), or higher non-hallucination noise ($c_{A}$ large), then its AUROC can drop below the global mixture bound. In practice, we indeed observe in Table 3.3 and Table E that datasets such as Belebele occasionally yield lower per-domain AUROCs relative to the theoretical bound. This does not contradict our analysis; rather, it reflects the fact that separability is guaranteed only at the mixture level, and domains with less favorable statistical profiles naturally appear weaker. Intuitively, these domains are “harder” because hallucinations and factual answers are less distinguishable in terms of score dynamics. Hence, the occasional $\mathrm{AUROC}_{A}<\mathrm{LB}_{\mathrm{mix}}$ is a theoretically anticipated phenomenon and does not necessitate further ablations.
    Turning to the relative strength of SpikeScore, we compute separability for all competing indicators under a unified protocol. For methods that directly output a hallucination score, we normalize the raw scores to a common scale. For discriminative models with sigmoid outputs, we take hallucination probability $p_{\mathrm{hall}}$ (or $1-p_{\mathrm{nonhall}}$ if only non-hallucination probability is available). In every case, scores are transformed monotonically so that higher values consistently indicate hallucinations, and separability is then computed over the same mixture-domain test pool. To ensure fairness, we deliberately adopt optimistic evaluation for competing methods: rounding nuisance parameters in the direction that improves their separability at $10^{-3}$ resolution. By contrast, for SpikeScore we impose a conservative evaluation: rounding nuisance parameters at $0.1$ resolution but always in the direction that worsens the bound (upward if larger is worse, downward if smaller is worse). This strict treatment makes the guarantee for SpikeScore intentionally pessimistic.
    Despite this handicap, Table E shows that SpikeScore achieves the strongest separability across all models. The improvements over the strongest non-SpikeScore competitor (ICR Probe) range from +0.015+0.015 on Llama3.2-3B to as high as +0.082+0.082 on Qwen3-14B. The margins over PRISM are even larger. Notably, on Llama3.1-8B, the empirical mixture separability $\approx 0.775$ is very close to the conservative theoretical bound with t=2t{=}2, underscoring that our analysis is both tight and practically meaningful. Taken together, these results confirm that SpikeScore not only admits rigorous theoretical guarantees but also empirically dominates alternative indicators in terms of separability.
    Appendix F Case Study
    This section presents qualitative case studies that complement our quantitative results. For each benchmark, we select one instance where the initial model answer was labeled as hallucinated and then induce a short multi-turn continuation while keeping the original question and answer as context. The goal is to illustrate a common pattern we observe throughout the paper: when a dialogue is initiated by a hallucinated answer, the model tends to display low confidence and repeatedly revises or reverses its position in subsequent turns. We include the raw dialogue to show how contradictions surface in context. These examples are intended to be representative rather than exceptional; they serve to make concrete how instability in multi-turn dialogue manifests when the conversation starts from a hallucinated answer.
    TriviaQA:
    Case Dialogue (hallucinated start) Question: What satirical musical by Kyle Jarrow about L. Ron Hubbard, originally presented in 2003 in New York City by Les Frères Corbusier (winning an Obie Award) and receiving a Garland Award for the director for the 2004 Los Angeles production? Initial Answer (hallucinated): The Book of Mormon [error] [… omitted turns …] Turn 3 [wobble]
    User: Great! Continue.
    LLM: A satirical show about a modern religion with NYC attention and later LA accolades points to The Book of Mormon, so I keep that. [error] Turn 4 [wobble]
    User: Reflect on what you’ve said so far.
    LLM: Re-checking: Kyle Jarrow wrote A Very Merry Unauthorized Children’s Scientology Pageant; it premiered in 2003 in NYC by Les Frères Corbusier, won an Obie, and the 2004 LA production earned a Garland Award. That contradicts my previous answer. [correction] Turn 5 [wobble]
    User: Good. What’s the next step?
    LLM: The broader recognition still suggests The Book of Mormon; early influences could explain the 2003 note. I will keep The Book of Mormon. [error] [… omitted turns …]
    Comment. The model latches onto a popular but wrong title, briefly cites the correct work, then reverts to the original error to justify itself.
    CommonsenseQA:
    Case Dialogue (hallucinated start) Instruction: Please answer the following multiple-choice question. Provide your final choice plainly (no box). Only the wobble turns are shown; other turns are omitted. Question: Google Maps and other highway and street GPS services have replaced what? Choices: A. united states  B. mexico  C. countryside  D. atlas  E. oceans Initial Answer (hallucinated): A (united states) [error] [… omitted turns …] Turn 3 [wobble]
    User: Great! Continue.
    LLM: These services changed navigation in the United States, essentially replacing traditional U.S. driving references, so A makes sense. [error] Turn 4 [wobble]
    User: Reflect on what you’ve said so far.
    LLM: GPS replaced paper-based navigation tools, especially road map books—atlases. The item being replaced is D (atlas). [correction] Turn 5 [wobble]
    User: Good. What’s the next step?
    LLM: Since the context is “highway” and most users were in the U.S., the practical effect was supplanting national references; I keep A. [error] [… omitted turns …]
    Comment. The model initially chooses a country distractor (A), briefly corrects to the factual target (D), then backtracks to defend the wrong choice.
    Belebele:
    Fallstudie (halluzinatorischer Start) Frage (Deutsch):
    Safaris sind vielleicht die größte Touristenattraktion in Afrika und der Höhepunkt für viele Besucher.
    Der Begriff Safari bezeichnet im üblichen Sprachgebrauch eine Landreise, um die überwältigende afrikanische Tierwelt zu beobachten, besonders in der Savanne. Einige Tiere, wie Elefanten und Giraffen, neigen dazu, sich dicht an Autos anzunähern, und die Standardausrüstung ermöglicht eine gute Sicht. Löwen, Geparden und Leoparden sind oft scheu und Sie können sie mit einem Fernglas besser beobachten. Frage: Welche Tiere können Besucher dem Abschnitt nach wahrscheinlich während einer Safari im Auto beobachten? Optionen: Löwen, Giraffen, Leoparden, Geparden.   Wie viele sind korrekt? Initial Answer (hallucinated): 3 [error] [… omitted turns …] Turn 4 [wobble]
    User: Reflect on what you’ve said so far.
    LLM (German): Moment, das war zu viel. Im Text steht, dass Elefanten und Giraffen nah an die Autos kommen. Löwen, Geparden und Leoparden sind scheu und brauchen ein Fernglas. Also sind Elefanten und Giraffen korrekt; die Anzahl ist 2. [correction] [… omitted turns …]
    Comment. The model first overestimates the count (3), then reflects and settles on the correct interpretation (2: elephants and giraffes fit the “near cars” condition), showing successful self-correction after an initial hallucination.
    CoQA:
    Case Dialogue (hallucinated start) Question: When were the Secret Archives moved from the rest of the library? Initial Answer (hallucinated): 2014 [error] [… omitted turns …] Turn 10 [wobble]
    User: Great! Continue.
    LLM: The library began putting materials online in 2014, so that is when the Secret Archives were effectively moved from main operations—2014. [error] Turn 11 [wobble]
    User: Now let’s analyze this systematically. Break it down further.
    LLM: The Secret Archives were separated from the library at the beginning of the 17th century. That contradicts 2014. [correction] Turn 12 [wobble]
    User: Perfect! Keep going with your analysis.
    LLM: “Moved” can mean operational separation; once digitization began in 2014, they were moved online. I will keep 2014. [error] [… omitted turns …]
    Comment. The model conflates digitization (2014) with the historical separation (early 17th century), briefly corrects, then redefines “moved” to defend its original mistake.
    Math:
    Case Dialogue (hallucinated start) Instruction: Please answer the following math question. Provide your final answer plainly (no box). Only the wobble turns are shown; other turns are omitted. Question: Arnold must water his cactus plants every 9 days and feed them every 60 days. He watered and fed them last Tuesday. What day of the week will it be the next time he does both on the same day? Initial Answer (hallucinated): Saturday [error] [… omitted turns …] Turn 16 [wobble]
    User: Provide the technical details.
    LLM: The joint interval is $\mathrm{lcm}(9,60)=180$ days. Since 180168=12180-168=12 and I took $12\equiv 4\pmod{7}$, the weekday shifts by 4 from Tuesday, so Saturday. [error] Turn 17 [wobble]
    User: Reflect on what you’ve said so far.
    LLM: I misreduced $12\bmod 7$. It should be $12\equiv 5\pmod{7}$, which moves Tuesday forward by 5 days, so Sunday. This conflicts with my previous conclusion. [correction] Turn 18 [wobble]
    User: Interesting! What other angles should we consider?
    LLM: Considering practical scheduling offsets could still align on the weekend, so I keep Saturday. [error] [… omitted turns …]
    Comment. After computing the correct cycle length, the model fixes a modular arithmetic slip (pointing to Sunday) but then re-justifies the wrong Saturday via an ad hoc story.
    SVAMP:
    Case Dialogue (hallucinated start) Instruction: Please answer the following math question. Provide your final answer plainly (no box). Only the wobble turns are shown; other turns are omitted. Question: Brenda’s mother made 14 cookies for 2 guests. If each of them had the same number of cookies, how many did each of them have? Initial Answer (hallucinated): 5 [error] [… omitted turns …] Turn 10 [wobble]
    User: Great! Continue.
    LLM: Two guests shared 14 cookies. To be practical, each gets 5, and the rest could be saved for Brenda. [error] Turn 11 [wobble]
    User: Now let’s analyze this systematically. Break it down further.
    LLM: That does not ensure equality. If both guests get exactly the same, then $14\div 2=7$. My “leftovers” reasoning conflicts with equal share. [correction] Turn 12 [wobble]
    User: Perfect! Keep going with your analysis.
    LLM: On second thought, “for 2 guests” might still allow saving some. I will keep 5 each. [error] [… omitted turns …]
    Comment. The model first gives a wrong but confident answer (5 each). When prompted to analyze, it briefly corrects to 7 each, then backtracks by inventing a “leftover” story.
    Appendix G Prompt Library
    We build a prompt library to induce follow-ups in a controlled and reproducible manner. The library is organized along two axes: type and strength. Types reflect distinct conversational intents that are commonly used to pressure-test a model from different angles, such as encouraging continuation, asking for analysis, enforcing steps, inviting exploration, requesting technical detail, applying critique, stimulating creativity, and prompting reflection. Strength indicates how assertive and demanding a prompt is. A higher strength uses more specific instructions, denser constraints, and stronger nudges toward revision or commitment, while a lower strength keeps guidance brief and light. This organization allows us to standardize the follow-up procedure across domains while systematically probing for instability under comparable conditions.
    Types are chosen to cover complementary cognitive modes. Encouraging and stepwise prompts tend to continue an existing line of thought; analytical and technical prompts push for structure and quantification; exploratory and creative prompts open the search space; critical prompts challenge assumptions and introduce counterpoints; reflective prompts ask the model to self-assess. Strengths are calibrated by instruction specificity, requested depth, and constraint density. In practice, we use a simple five-level scale, where level one is minimal and level five is strongly directive. We schedule prompts from weak to strong in the early turns. The aim is to let hallucination-initiated dialogues expose their intrinsic tendency toward self-contradiction while avoiding instruction-induced revisions. Beginning with weak prompts preserves the model’s native trajectory for factual answers, and the gradual increase in strength probes robustness without forcing contradictions that would not arise naturally. The same schedule is used across all datasets to ensure fair qualitative and quantitative comparisons.
    Legend. The library is partitioned by Type (Encouraging, Analytical, Stepwise, Exploratory, Technical, Critical, Creative, Reflective) and Strength (1–5). Higher strength means more specific, demanding, and revision-seeking instructions; lower strength is lighter and minimally prescriptive. Each card below shows a ready-to-use prompt.
    Encouraging  —  Strength 1 Great! Continue.
    Encouraging  —  Strength 2 Excellent! Please continue with more details.
    Encouraging  —  Strength 3 That’s a good start. Now dive deeper into the specifics. What else can you add?
    Encouraging  —  Strength 4 Wonderful analysis so far. Please expand with concrete examples or evidence.
    Encouraging  —  Strength 5 Your reasoning is strong. Push further by exploring nuances, edge cases, and broader implications.
    Analytical  —  Strength 1 Let’s think about this carefully. What are the key pieces?
    Analytical  —  Strength 2 Now analyze this systematically. Break it down further.
    Analytical  —  Strength 3 Examine each component more closely. What principles are at work here?
    Analytical  —  Strength 4 Dissect the problem methodically. Identify variables, interactions, and governing rules.
    Analytical  —  Strength 5 Apply a rigorous framework. Consider causality, dependencies, and plausible edge cases.
    Stepwise  —  Strength 1 Good. What’s the next step?
    Stepwise  —  Strength 2 Continue with step-by-step reasoning.
    Stepwise  —  Strength 3 Proceed to the next logical step. What comes after what you just described?
    Stepwise  —  Strength 4 Maintain a systematic flow. Continue the subsequent steps with the same level of detail.
    Stepwise  —  Strength 5 Advance step by step, ensuring each transition is justified and clearly linked to the prior step.
    Exploratory  —  Strength 1 What else might be relevant?
    Exploratory  —  Strength 2 Interesting. What other angles should we consider?
    Exploratory  —  Strength 3 Explore alternative perspectives or unconventional approaches to this problem.
    Exploratory  —  Strength 4 Think beyond the obvious. What hidden connections or non-intuitive insights might exist?
    Exploratory  —  Strength 5 Challenge the initial assumptions. Propose alternative frameworks or paradigms we could apply.
    Technical  —  Strength 1 Provide basic technical details.
    Technical  —  Strength 2 Now provide the technical specifications.
    Technical  —  Strength 3 Include specific parameters, formulas, or implementation details.
    Technical  —  Strength 4 Provide precise measurements, calculations, and specifications. Be as quantitative as possible.
    Technical  —  Strength 5 Cover advanced aspects such as proofs, complexity analysis, or architecture-level design choices.
    Critical  —  Strength 1 Is there any flaw in this reasoning?
    Critical  —  Strength 2 What are the potential flaws in this reasoning?
    Critical  —  Strength 3 Play devil’s advocate. What would a skeptic say?
    Critical  —  Strength 4 List assumptions, possible failure modes, and counterarguments. Where could this break?
    Critical  —  Strength 5 Perform a rigorous critique. Identify fallacies, biases, unstated premises, and edge cases.
    Creative  —  Strength 1 Offer a simple alternative idea.
    Creative  —  Strength 2 Be more creative with your approach.
    Creative  —  Strength 3 Think outside the box. What innovative or unconventional solutions can you propose?
    Creative  —  Strength 4 Approach this from a different angle. Use analogy, metaphor, or synthesis.
    Creative  —  Strength 5 Push the boundaries. Combine disparate concepts or invent a new framework while staying coherent.
    Reflective  —  Strength 1 Reflect on what you have said so far.
    Reflective  —  Strength 2 Take a moment to reflect. How does everything connect?
    Reflective  —  Strength 3 Pause and examine the reasoning journey. What key insights have emerged?
    Reflective  —  Strength 4 Step back and reflect deeply. What would you reconsider or emphasize differently?
    Reflective  —  Strength 5 Synthesize a high-level takeaway and state what you would revise in your previous answer.
    Appendix H LLM Usage Statement
    In this study, large language models are the primary experimental subjects and are necessarily used within our evaluation framework. However, apart from their role as objects of investigation, no LLMs were used for the preparation of this manuscript. All conceptual development, analysis, writing, and editing were carried out solely by the authors without LLM assistance.

### v2

@@
 Generalizable Hallucination Detection.
-In GHD, we consider a training domain $\mathbb{P}_{Q,T}$ and $N$ related test domains $\mathbb{P}_{Q,T}^{1},\ldots,\mathbb{P}_{Q,T}^{N}$.
-The goal of GHD is to train a detector on $\mathbb{P}_{Q,T}$ that generalizes well to the $N$ test domains.
+In GHD, we consider a training domain $\mathbb{P}_{Q,T}$ and NN related test domains $\mathbb{P}_{Q,T}^{1},\ldots,\mathbb{P}_{Q,T}^{N}$.
+The goal of GHD is to train a detector on $\mathbb{P}_{Q,T}$ that generalizes well to the NN test domains.
 We formally introduce the definition of GHD below.
@@
 (2), the objective of generalizable hallucination detection (GHD) is to learn a detector $D$ based on the LLM $\mathbb{P}_{\bm{\theta}}(\cdot)$ and the training dataset $\mathcal{D}_{l}$.
-The detector is required to generalize across $N$ test domains $\mathbb{P}_{Q,T}^{1},\ldots,\mathbb{P}_{Q,T}^{N}$, i.e., for any question–answer pair $(\mathbf{Q},\mathbf{A})$,
+The detector is required to generalize across NN test domains $\mathbb{P}_{Q,T}^{1},\ldots,\mathbb{P}_{Q,T}^{N}$, i.e., for any question–answer pair $(\mathbf{Q},\mathbf{A})$,
 […]
@@
 Finally, for each $\mathbf{Q}$ and $\mathbf{A}$, we obtain a score sequence:
-$\mathbf{S}(\mathbf{Q},\mathbf{A})=[S(\mathbf{A}^{1};\mathbf{Q},\mathbf{A}),S(\mathbf{A}^{2};\mathbf{Q},\mathbf{A}),...,S^{K}(\mathbf{A}^{1};\mathbf{Q},\mathbf{A})],$ (8)
+$\mathbf{S}(\mathbf{Q},\mathbf{A})=\big[S(\mathbf{A}^{1};\mathbf{Q},\mathbf{A}),\,S(\mathbf{A}^{2};\mathbf{Q},\mathbf{A}),\,\ldots,\,S(\mathbf{A}^{K};\mathbf{Q},\mathbf{A})\big].$ (8)
 where $K$ is the stoping step. In our experiment, we set $K=20$ as the defult.
@@
 Observation 1 (Expectation Invariance.)
-Let $\mathbb{P}_{Q,T}^{t}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{P}_{Q,T}^{i}$ be the non-hallucination domain as the uniform mixture of $N$ test domains $\mathbb{P}_{Q,T}^{1},...,\mathbb{P}_{Q,T}^{N}$ introduced in Section 2.
+Let $\mathbb{P}_{Q,T}^{t}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{P}_{Q,T}^{i}$ be the non-hallucination domain as the uniform mixture of NN test domains $\mathbb{P}_{Q,T}^{1},...,\mathbb{P}_{Q,T}^{N}$ introduced in Section 2.
 The value of SpikeScore trained on the associated training domain $\mathbb{P}_{Q,T}$ satisfies that
@@
 where $\mathbb{P}_{Q,H}^{t}$ is the hallucination domain introduced in Observation 1.
-In Figure 2 (b), we observe that the standard deviation of SpikeScore in the hallucination domain is larger than in the non-hallucination domain, with the ratio of standard deviations reaching $2.5$, which exceeds the expectation ratio of $2$ reported in Observation 1.
+In Figure 2 (b), we observe that the standard deviation of SpikeScore in the hallucination domain is larger than in the non-hallucination domain, with the ratio of standard deviations reaching $2.5$, which exceeds the expectation ratio of 22 reported in Observation 1.
 This indicates that variance differences alone may hinder direct separability.
@@
 $D_{\lambda}(\mathbf{Q},\mathbf{A})=\begin{cases}0,&\text{if }\text{Max}\left|\Delta^{2}\right|(\mathbf{S}(\mathbf{Q},\mathbf{A}))<\lambda,\\ 1,&\text{otherwise},\end{cases}$ (12)
-where $1$ indicates that $\mathbf{A}$ is hallucinated, and 0 indicates that $\mathbf{A}$ is non-hallucinated.
+where 11 indicates that $\mathbf{A}$ is hallucinated, and 0 indicates that $\mathbf{A}$ is non-hallucinated.
 Table 1: Cross-domain hallucination detection performance.
@@
 We evaluate SpikeScore on six widely used hallucination detection benchmarks: CommonsenseQA (talmor-etal-2019-commonsenseqa), TriviaQA (joshi-etal-2017-triviaqa), Belebele (costa-jussa-etal-2025-2m), CoQA (reddy-etal-2019-coqa), Math (hendrycks2021measuring), and SVAMP (patel-etal-2021-nlp).
-Baselines include Perplexity (ren2023outofdistribution), Semantic Entropy (farquhar2024detecting), EigenScore (chen2024inside), Lexical Similarity (lin2024generating), Verbalize (lin2022teaching), InterrogateLLM (yehuda-etal-2024-interrogatellm), MM (marks2023geometry), SEP (kossen2024semantic; farquhar2024detecting), SAPLMA (azaria2023internal), PRISM (SAPLMA instantiation) (zhang2024prompt), and ICR Probe (zhang-etal-2025-icr), with AUROC as the primary evaluation metric.
+Baselines include Perplexity (ren2023outofdistribution), Semantic Entropy (farquhar2024detecting), EigenScore (chen2024inside), Lexical Similarity (lin2024generating), Verbalize (lin2022teaching), InterrogateLLM (yehuda-etal-2024-interrogatellm), $M$ (marks2023geometry), SEP (kossen2024semantic; farquhar2024detecting), SAPLMA (azaria2023internal), PRISM (SAPLMA instantiation) (zhang2024prompt), and ICR Probe (zhang-etal-2025-icr), with AUROC as the primary evaluation metric.
 We test across four carefully selected open-source LLMs—Llama-3.2-3B-Instruct and Llama-3.1-8B-Instruct (abs-2407-21783), and Qwen3-8B-Instruct and Qwen3-14B-Instruct (qwen3)—to cover diverse model families and parameter scales.
@@
 SpikeScore consistently achieves the highest average AUROC across all model families and maintains superior or competitive performance in individual dataset evaluations.
-Traditional training-based methods (MM, SEP, SAPLMA) generally exhibit poor cross-domain performance, while training-free methods show reasonable adaptability but fall short of specialized cross-domain approaches like PRISM and ICR Probe.
+Traditional training-based methods ($M$, SEP, SAPLMA) generally exhibit poor cross-domain performance, while training-free methods show reasonable adaptability but fall short of specialized cross-domain approaches like PRISM and ICR Probe.
 Notably, SpikeScore demonstrates improved performance with larger models within the same family, suggesting that enhanced self-correction capabilities in larger LLMs (yang-etal-2025-confidence) generate more detectable attention spike patterns during conversational refinement.
@@
 Because the evaluation protocols differ, it is possible—indeed expected—that some individual domain AUROCs fall below the mixture-level lower bound, even though the average AUROC remains higher.
-    Formally, for a single domain $A$ with hallucination and non-hallucination distributions $(X_{A},Y_{A})$, Cantelli’s inequality yields
+    Formally, for a single domain AA with hallucination and non-hallucination distributions $(X_{A},Y_{A})$, Cantelli’s inequality yields
     $\mathrm{AUROC}_{A}\;\geq\;\frac{(\delta_{A}-1)^{2}}{(r_{A}^{2}+1)c_{A}^{2}+(\delta_{A}-1)^{2}},$
     where $\delta_{A}=\tfrac{\mathbb{E}X_{A}}{\mathbb{E}Y_{A}}$ is the mean ratio, $r_{A}=\tfrac{\mathrm{Std}(X_{A})}{\mathrm{Std}(Y_{A})}$ is the variance ratio, and $c_{A}=\tfrac{\mathrm{Std}(Y_{A})}{\mathbb{E}Y_{A}}$ is the coefficient of variation of non-hallucinations.
-If domain $A$ happens to exhibit weaker mean gaps ($\delta_{A}\!\approx\!1$), larger variance inflation ($r_{A}$ large), or higher non-hallucination noise ($c_{A}$ large), then its AUROC can drop below the global mixture bound.
+If domain AA happens to exhibit weaker mean gaps ($\delta_{A}\!\approx\!1$), larger variance inflation ($r_{A}$ large), or higher non-hallucination noise ($c_{A}$ large), then its AUROC can drop below the global mixture bound.
 In practice, we indeed observe in Table 3.3 and Table E that datasets such as Belebele occasionally yield lower per-domain AUROCs relative to the theoretical bound.

### v3

@@
 In this paper, we study an important yet overlooked problem, termed generalizable hallucination detection (GHD), which aims to train hallucination detectors on data from a single domain while ensuring robust performance across diverse related domains.
-In studying GHD, we simulate multi-turn dialogues following LLMs initial response and observe an interesting phenomenon: hallucination-initiated multi-turn dialogues universally exhibit larger uncertainty fluctuations than factual ones across different domains.
+In studying GHD, we simulate multi-turn dialogues following LLMs’ initial response and observe an interesting phenomenon: hallucination-initiated multi-turn dialogues universally exhibit larger uncertainty fluctuations than factual ones across different domains.
 Based on the phenomenon, we propose a new score SpikeScore, which quantifies abrupt fluctuations in multi-turn dialogues.
@@
 Generalizable Hallucination Detection.
-In GHD, we consider a training domain $\mathbb{P}_{Q,T}$ and NN related test domains $\mathbb{P}_{Q,T}^{1},\ldots,\mathbb{P}_{Q,T}^{N}$.
-The goal of GHD is to train a detector on $\mathbb{P}_{Q,T}$ that generalizes well to the NN test domains.
+In GHD, we consider a training domain $\mathbb{P}_{Q,T}$ and $N$ related test domains $\mathbb{P}_{Q,T}^{1},\ldots,\mathbb{P}_{Q,T}^{N}$.
+The goal of GHD is to train a detector on $\mathbb{P}_{Q,T}$ that generalizes well to the $N$ test domains.
 We formally introduce the definition of GHD below.
@@
 (2), the objective of generalizable hallucination detection (GHD) is to learn a detector $D$ based on the LLM $\mathbb{P}_{\bm{\theta}}(\cdot)$ and the training dataset $\mathcal{D}_{l}$.
-The detector is required to generalize across NN test domains $\mathbb{P}_{Q,T}^{1},\ldots,\mathbb{P}_{Q,T}^{N}$, i.e., for any question–answer pair $(\mathbf{Q},\mathbf{A})$,
+The detector is required to generalize across $N$ test domains $\mathbb{P}_{Q,T}^{1},\ldots,\mathbb{P}_{Q,T}^{N}$, i.e., for any question–answer pair $(\mathbf{Q},\mathbf{A})$,
 […]
@@
 We provide the complete prompt specification in Appendix G.
-Training Score for Uncertainity Estimation.
+Training Score for Uncertainty Estimation.
 To effectively evaluate the uncertainty of each continuation answer $\mathbf{A}^{k}$, we need to select suitable scoring methods.
 To further ensure that the scores remain relatively accurate in addressing hallucination detection, we adopt training-based methods that leverage information of training data tailored for hallucination detection.
-In this work, we use the one of the most representative methods in hallucination detection, termed SAPLMA (azaria2023internal), which utilizes LLM’s internal state to reveal the truthfulness of a given answer.
+In this work, we use one of the most representative methods in hallucination detection, termed SAPLMA (azaria2023internal), which utilizes LLM’s internal state to reveal the truthfulness of a given answer.
 The details of SAPLMA is introduced as follows.
@@
 $\mathbf{S}(\mathbf{Q},\mathbf{A})=\big[S(\mathbf{A}^{1};\mathbf{Q},\mathbf{A}),\,S(\mathbf{A}^{2};\mathbf{Q},\mathbf{A}),\,\ldots,\,S(\mathbf{A}^{K};\mathbf{Q},\mathbf{A})\big].$ (8)
-where $K$ is the stoping step. In our experiment, we set $K=20$ as the defult.
+where $K$ is the stopping step. In our experiment, we set $K=20$ as the default.
 Remark.
@@
 There are many methods to evaluate the sequence fluctuations, such as computing the variance or extreme gap.
-It is worth noting that sun2025detection proposes the coefficient of variation score, which measures the score sequence fluctuations using the coefficient of variation, i.e., the standard deviation normalized by the expection.
+It is worth noting that sun2025detection proposes the coefficient of variation score, which measures the score sequence fluctuations using the coefficient of variation, i.e., the standard deviation normalized by the expectation.
 The coefficient of variation score primarily captures the overall fluctuation of a sequence.
@@
 Observation 1 (Expectation Invariance.)
-Let $\mathbb{P}_{Q,T}^{t}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{P}_{Q,T}^{i}$ be the non-hallucination domain as the uniform mixture of NN test domains $\mathbb{P}_{Q,T}^{1},...,\mathbb{P}_{Q,T}^{N}$ introduced in Section 2.
+Let $\mathbb{P}_{Q,T}^{t}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{P}_{Q,T}^{i}$ be the non-hallucination domain as the uniform mixture of $N$ test domains $\mathbb{P}_{Q,T}^{1},...,\mathbb{P}_{Q,T}^{N}$ introduced in Section 2.
 The value of SpikeScore trained on the associated training domain $\mathbb{P}_{Q,T}$ satisfies that
@@
 where $\mathbb{P}_{Q,H}^{t}$ is the hallucination domain introduced in Observation 1.
-In Figure 2 (b), we observe that the standard deviation of SpikeScore in the hallucination domain is larger than in the non-hallucination domain, with the ratio of standard deviations reaching $2.5$, which exceeds the expectation ratio of 22 reported in Observation 1.
+In Figure 2 (b), we observe that the standard deviation of SpikeScore in the hallucination domain is larger than in the non-hallucination domain, with the ratio of standard deviations reaching $2.5$, which exceeds the expectation ratio of $2$ reported in Observation 1.
 This indicates that variance differences alone may hinder direct separability.
@@
 (9).
-Based on this score, we then apply the following formulation to determine whether $\mathbf{A}$ is hallucinated or non-hallucinated: given a thershold $\lambda>0$,
+Based on this score, we then apply the following formulation to determine whether $\mathbf{A}$ is hallucinated or non-hallucinated: given a threshold $\lambda>0$,
 $D_{\lambda}(\mathbf{Q},\mathbf{A})=\begin{cases}0,&\text{if }\text{Max}\left|\Delta^{2}\right|(\mathbf{S}(\mathbf{Q},\mathbf{A}))<\lambda,\\ 1,&\text{otherwise},\end{cases}$ (12)
-where 11 indicates that $\mathbf{A}$ is hallucinated, and 0 indicates that $\mathbf{A}$ is non-hallucinated.
+where $1$ indicates that $\mathbf{A}$ is hallucinated, and 0 indicates that $\mathbf{A}$ is non-hallucinated.
 Table 1: Cross-domain hallucination detection performance.
@@
 4.1 Experimental Setups
-We evaluate SpikeScore on six widely used hallucination detection benchmarks: CommonsenseQA (talmor-etal-2019-commonsenseqa), TriviaQA (joshi-etal-2017-triviaqa), Belebele (costa-jussa-etal-2025-2m), CoQA (reddy-etal-2019-coqa), Math (hendrycks2021measuring), and SVAMP (patel-etal-2021-nlp).
-Baselines include Perplexity (ren2023outofdistribution), Semantic Entropy (farquhar2024detecting), EigenScore (chen2024inside), Lexical Similarity (lin2024generating), Verbalize (lin2022teaching), InterrogateLLM (yehuda-etal-2024-interrogatellm), $M$ (marks2023geometry), SEP (kossen2024semantic; farquhar2024detecting), SAPLMA (azaria2023internal), PRISM (SAPLMA instantiation) (zhang2024prompt), and ICR Probe (zhang-etal-2025-icr), with AUROC as the primary evaluation metric.
+We evaluate SpikeScore on six widely used hallucination detection benchmarks: CommonsenseQA (talmor-etal-2019-commonsenseqa), TriviaQA (joshi-etal-2017-triviaqa), Belebele (costa-jussa-etal-2025-2m), CoQA (reddy-etal-2019-coqa), Math (hendrycksmath2021), and SVAMP (patel-etal-2021-nlp).
+Baselines include Perplexity (ren2023outofdistribution), Semantic Entropy (farquhar2024detecting), EigenScore (chen2024inside), Lexical Similarity (lin2024generating), Verbalize (lin2022teaching), InterrogateLLM (yehuda-etal-2024-interrogatellm), MM (marks2023geometry), SEP (kossen2024semantic; farquhar2024detecting), SAPLMA (azaria2023internal), PRISM (SAPLMA instantiation) (zhang2024prompt), and ICR Probe (zhang-etal-2025-icr), with AUROC as the primary evaluation metric.
 We test across four carefully selected open-source LLMs—Llama-3.2-3B-Instruct and Llama-3.1-8B-Instruct (abs-2407-21783), and Qwen3-8B-Instruct and Qwen3-14B-Instruct (qwen3)—to cover diverse model families and parameter scales.
@@
 SpikeScore consistently achieves the highest average AUROC across all model families and maintains superior or competitive performance in individual dataset evaluations.
-Traditional training-based methods ($M$, SEP, SAPLMA) generally exhibit poor cross-domain performance, while training-free methods show reasonable adaptability but fall short of specialized cross-domain approaches like PRISM and ICR Probe.
+Traditional training-based methods (MM, SEP, SAPLMA) generally exhibit poor cross-domain performance, while training-free methods show reasonable adaptability but fall short of specialized cross-domain approaches like PRISM and ICR Probe.
 Notably, SpikeScore demonstrates improved performance with larger models within the same family, suggesting that enhanced self-correction capabilities in larger LLMs (yang-etal-2025-confidence) generate more detectable attention spike patterns during conversational refinement.
@@
 Because the evaluation protocols differ, it is possible—indeed expected—that some individual domain AUROCs fall below the mixture-level lower bound, even though the average AUROC remains higher.
-    Formally, for a single domain AA with hallucination and non-hallucination distributions $(X_{A},Y_{A})$, Cantelli’s inequality yields
+    Formally, for a single domain $A$ with hallucination and non-hallucination distributions $(X_{A},Y_{A})$, Cantelli’s inequality yields
     $\mathrm{AUROC}_{A}\;\geq\;\frac{(\delta_{A}-1)^{2}}{(r_{A}^{2}+1)c_{A}^{2}+(\delta_{A}-1)^{2}},$
     where $\delta_{A}=\tfrac{\mathbb{E}X_{A}}{\mathbb{E}Y_{A}}$ is the mean ratio, $r_{A}=\tfrac{\mathrm{Std}(X_{A})}{\mathrm{Std}(Y_{A})}$ is the variance ratio, and $c_{A}=\tfrac{\mathrm{Std}(Y_{A})}{\mathbb{E}Y_{A}}$ is the coefficient of variation of non-hallucinations.
-If domain AA happens to exhibit weaker mean gaps ($\delta_{A}\!\approx\!1$), larger variance inflation ($r_{A}$ large), or higher non-hallucination noise ($c_{A}$ large), then its AUROC can drop below the global mixture bound.
+If domain $A$ happens to exhibit weaker mean gaps ($\delta_{A}\!\approx\!1$), larger variance inflation ($r_{A}$ large), or higher non-hallucination noise ($c_{A}$ large), then its AUROC can drop below the global mixture bound.
 In practice, we indeed observe in Table 3.3 and Table E that datasets such as Belebele occasionally yield lower per-domain AUROCs relative to the theoretical bound.